# SUDDEN CHANGE OF INTERFEROMETRIC POWER FOR X SHAPE STATES

### v1

@@ 2 Quantum IP as the discordlike measure of X shape states
One can note that IP has the following properties: (i) $\mathcal{IP}(\rho_{AB})$ equals to zero if $\rho_{AB}$ is a classical state (with respect to $A$); (ii) $\mathcal{IP}(\rho_{AB})$ is invariant under local unitary operation; (iii) $\mathcal{IP}(\rho_{AB})$ is monotonically decreasing under local completely positive and trace preserving maps on subsystem $B$; (iv) $\mathcal{IP}(\rho_{AB})$ reduces to a measure of entanglement if $\rho_{AB}$ is a pure state.[29] These properties imply that $\mathcal{IP}(\rho_{AB})$ is a proper measure of discord-type correlation.
@@ 3.1 Amplitude noise
$\begin{split}K_{1a}&=\left(\begin{array}[]{cc}\eta&0\\ 0&1\end{array}\right)\otimes\left(\begin{array}[]{cc}1&0\\ 0&1\end{array}\right),\\ K_{2a}&=\left(\begin{array}[]{cc}0&0\\ \sqrt{1-\eta^{2}}&1\end{array}\right)\otimes\left(\begin{array}[]{cc}1&0\\ 0&1\end{array}\right),\\ \end{split}$ (10)
@@ 3.2 Phase noise
$\begin{split}K_{1p}&=\sqrt{\alpha}\left(\begin{array}[]{cc}1&0\\ 0&1\end{array}\right)\otimes\left(\begin{array}[]{cc}1&0\\ 0&1\end{array}\right),\\ K_{2p}&=\sqrt{1-\alpha}\left(\begin{array}[]{cc}1&0\\ 0&-1\end{array}\right)\otimes\left(\begin{array}[]{cc}1&0\\ 0&1\end{array}\right),\\ \end{split}$ (14)
@@
where $\|C(t)\|^{2}=\mathsf{Tr}[C^{T}C]=c_{1}^{2}(t)+c_{2}^{2}(t)+c_{3}^{2}(t)$ the square Hilbert-Schmidt of $C(t)$ and $\|C(t)\|_{\infty}^{2}=\max\{c_{1}^{2}(t),c_{2}^{2}(t),c_{3}^{2}(t)\}$ the operator norm of $C(t)$.[29] If $|c_{3}|\geq\max\{|c_{1}|,|c_{2}|\}$, $\|C(t)\|_{\infty}^{2}$ will reduce to $c_{3}^{2}(t)$ since the decay rate of $\max\{|c_{1}(t)|,|c_{2}(t)|\}$ and $|c_{3}(t)|$ is different, then the IP $\mathcal{IP}(\mathcal{E}_{p}(\rho))$ decays monotonically. If $|c_{3}|<\max\{|c_{1}|,|c_{2}|\}$ and $c_{3}\neq 0$, the decay rate of IP $\mathcal{IP}(\mathcal{E}_{p}(\rho))$ sudden changes at $t_{0}=-\frac{2}{\tau}\ln\left|\frac{c_{3}}{\max\{|c_{1}|,|c_{2}|\}}\right|$. In Fig. 2, we choose three different $\{c_{i}\}$ to represent the dynamics of IP of $\rho$ under single qubit phase noise. As a result, a sudden change of IP may occur under this kind of noise environment.
@@ 3.3 Depolarizing noise
As one of the important types of quantum noise, the depolarizing noise describes a process that takes a state into completely mixed state $\mathbb{1}/2$ with probability $p$ and the state being left untouched with probability $1-p$. The operation elements for depolarizing noise are shown as $\{\sqrt{1-3p/4}\mathbb{1},\sqrt{p}\sigma_{x}/2,\sqrt{p}\sigma_{y}/2,\sqrt{p}\sigma_{z}/2\}$ and the Kraus operators for the whole system are given by[1]

### v2

@@ 2 Quantum IP as the discordlike measure of X shape states
One can note that IP has the following properties: (i) $\mathcal{IP}(\rho_{AB})$ equals to zero if $\rho_{AB}$ is a classical state (with respect to AA); (ii) $\mathcal{IP}(\rho_{AB})$ is invariant under local unitary operation; (iii) $\mathcal{IP}(\rho_{AB})$ is monotonically decreasing under local completely positive and trace preserving maps on subsystem $B$; (iv) $\mathcal{IP}(\rho_{AB})$ reduces to a measure of entanglement if $\rho_{AB}$ is a pure state.[29] These properties imply that $\mathcal{IP}(\rho_{AB})$ is a proper measure of discord-type correlation.
@@ 3.1 Amplitude noise
[…]
@@ 3.2 Phase noise
[…]
@@
where $\|C(t)\|^{2}=\mathsf{Tr}[C^{T}C]=c_{1}^{2}(t)+c_{2}^{2}(t)+c_{3}^{2}(t)$ the square Hilbert-Schmidt of $C(t)$ and $\|C(t)\|_{\infty}^{2}=\max\{c_{1}^{2}(t),c_{2}^{2}(t),c_{3}^{2}(t)\}$ the operator norm of C(t)C(t).[29] If $|c_{3}|\geq\max\{|c_{1}|,|c_{2}|\}$, $\|C(t)\|_{\infty}^{2}$ will reduce to $c_{3}^{2}(t)$ since the decay rate of $\max\{|c_{1}(t)|,|c_{2}(t)|\}$ and $|c_{3}(t)|$ is different, then the IP $\mathcal{IP}(\mathcal{E}_{p}(\rho))$ decays monotonically. If $|c_{3}|<\max\{|c_{1}|,|c_{2}|\}$ and $c_{3}\neq 0$, the decay rate of IP $\mathcal{IP}(\mathcal{E}_{p}(\rho))$ sudden changes at $t_{0}=-\frac{2}{\tau}\ln\left|\frac{c_{3}}{\max\{|c_{1}|,|c_{2}|\}}\right|$. In Fig. 2, we choose three different $\{c_{i}\}$ to represent the dynamics of IP of $\rho$ under single qubit phase noise. As a result, a sudden change of IP may occur under this kind of noise environment.
@@ 3.3 Depolarizing noise
As one of the important types of quantum noise, the depolarizing noise describes a process that takes a state into completely mixed state $\mathbb{1}/2$ with probability $p$ and the state being left untouched with probability 1p1-p. The operation elements for depolarizing noise are shown as $\{\sqrt{1-3p/4}\mathbb{1},\sqrt{p}\sigma_{x}/2,\sqrt{p}\sigma_{y}/2,\sqrt{p}\sigma_{z}/2\}$ and the Kraus operators for the whole system are given by[1]